OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11

OpenStudy (anonymous):

(1) y=2-3x (2)y= -11-7x/-5

OpenStudy (anonymous):

choices... –2 –3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system

OpenStudy (anonymous):

place any value for x!!!!!

OpenStudy (anonymous):

y=-2

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

ya its -2 !!!!